Write the net ionic equation that shows the reaction of 1.0 M ammonia, NH3, when dissolved in water. At room temperature the KB of the above solution is 1.8*10^-5. If the initial concentration of NH3 is 1.0 M, quantitatively define the pH of the above solution (show calculations using an I.C.E table and use the quadratic equation).
Answer #1

               NH3 + H2O ----------> NH4+ + OH-

    I          1                                 0           0

   C          -x                               +x           +x

   E       1-x                                 +x           +x

           Kb   =[NH4+][OH-]/[NH3]

           1.8*10^-5 n = x*x/1-x

            1.8*10^-5*(1-x) = x^2

               x   = 0.00423

           [OH-] = x= 0.00423M

           POH   = -log[OH-]

                     = -log0.00423

                       = 2.3736

            PH      = 14-POH

                       = 14-2.3736   = 11.6264